OpenStudy (anonymous):

hello i need help,hurry im doing a timed test 4a+7b=-5 8a+4c=36 6b+4c=-14

OpenStudy (anonymous):

Did you use the calculator provided on this website? is this algebra? what are you trying to do, factoring?

OpenStudy (anonymous):

its algebra 2! and no which calculater do i use...its says to solve

OpenStudy (anonymous):

the system of equations

OpenStudy (anonymous):

solve by substitution? be specific?

OpenStudy (anonymous):

it didnt say.. it just says solve system of equations

OpenStudy (anonymous):

if it's Solve by Substitution Method: c=1

OpenStudy (anonymous):

according to the calculator provided here.

OpenStudy (anonymous):

okay so i plug 1 into all the c's?

OpenStudy (anonymous):

a=-7b/4-5/4 b=2c/7-23/7 c=1 according to the calculator provided in this website.

OpenStudy (anonymous):

that's if you are trying to Solve by Substitution Method:

OpenStudy (anonymous):

your thebest

OpenStudy (anonymous):

is it correct?

OpenStudy (anonymous):

idk i will find out what about this one it sayswithout graphing classify the following system as independent dependent or inconsistent

OpenStudy (anonymous):

2x-y=3 6x-3y=0

OpenStudy (anonymous):

according to the calculator provided it is incosistent

OpenStudy (anonymous):

okay thanks what about this one it says solve the systems of equattions useing matrix

OpenStudy (anonymous):

x+3y=22 2x-y=9

OpenStudy (anonymous):

it says x=7, y=5

OpenStudy (anonymous):

just go to this website and solve for yourself http://www.freemathhelp.com/algebra-help.html scroll all the way down and you will find the calculator.

OpenStudy (anonymous):

i dont understand how to use it help me with one more please

OpenStudy (anonymous):

solve the system by elimination 4a+4b-c+=-6 8a+6b+c=-5 6z-4b-6c=-22

OpenStudy (anonymous):

on the first equation after c+ then what?

OpenStudy (anonymous):

ooops let me write it over its 4a+4b-c=-6

OpenStudy (anonymous):

and the last one after the 6 its and a,,im in a hurry :)

OpenStudy (anonymous):

you mean this on the last eq. 6z-4b-6a=-22

OpenStudy (anonymous):

nooo the last one was 6a-4b-6b=-22

OpenStudy (anonymous):

is it Solve by Addition/Elimination Method:?

OpenStudy (anonymous):

probably?

OpenStudy (anonymous):

if so then the calculator says can't be solve in addition method but on Solve by Gaussian Elimination: its (-11/6, 11/10, 46/15)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks what about this one: find the maximum valube of the objective function and the calues of x and y for which it occurs

OpenStudy (anonymous):

f=2x+y 3x+5y(less than or greater then sign) 45 2x+4y(less than or greater then sign) 32

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

helllo